OpenStudy (anonymous):

What is the equation of the line passing through (2, –3) and parallel to the line y = –6x – 1 in standard form?

OpenStudy (anonymous):

parallel means it has the same slope, which is -6. you can use the point slope formula, y-y1 = m(x-x1). plug in your points to y1 and x1 to end up with y+3=-6(x-2). for standard form, just bring the 3 over to the right side and multiply out the parentheses.

OpenStudy (anonymous):

thank you